# A Survey on Occupancy Perception for Autonomous Driving: The Information Fusion Perspective

### v1

@@ 3.3.1 General Pipeline
RGB images captured by cameras provide rich and dense semantic information but are sensitive to weather condition changes and lack precise geometric details. In contrast, point clouds from LiDAR or radar are robust to weather changes and excel at capturing scene geometry with accurate depth measurements. However, they only produce sparse features. Multi-modal occupancy perception can combine the advantages from multiple modalities, and mitigate the limitations of single-modal perception. Fig. 3(c) illustrates the general pipeline of multi-modal occupancy perception. Most multi-modal methods [10, 87, 11, 14] map 2D image features into 3D space and then fuse them with point cloud features. Moreover, incorporating 2D perspective-view features in the fusion process can further refine the representation [13]. The fused representation is processed by an optional refinement module and an occupancy head, such as 3D convolution or MLP, to generate the final 3D occupancy predictions. The optional refinement module [88] could be a combination of cross attention, self attention, and diffusion denoising [108].
@@ 3.4.1 Training with Strong Supervision
where $\left(V,\hat{V}\right)$ are the ground-truth and predicted semantic occupancy with $N_{C}$ categories. $\omega_{c}$ is a weight for a specific class $c$ according to the inverse of the class frequency. Moreover, some occupancy perception methods use other semantic losses, which are commonly utilized in semantic segmentation tasks, such as Lovasz-Softmax loss [113] and Focal loss [114]. Furthermore, there are two specialized semantic losses: frustum proportion loss [23], which provides cues to alleviate occlusion ambiguities from the perspective of the visual frustum, and position awareness loss [115], which leverages local semantic entropy to encourage sharper semantic and geometric gradients.
@@ 4.1.1 Datasets
There are a variety of datasets to evaluate the performance of occupancy prediction approaches, e.g., the widely used KITTI [122], nuScenes [80], and Waymo [123]. However, most of the datasets only contain 2D semantic segmentation annotations, which is not practical for the training or evaluation of 3D occupancy prediction approaches. To support the benchmarks for 3D occupancy perception, many new datasets such as Monoscene [23], Occ3D [33], and OpenScene [127] are developed based on the previous datasets like nuScenes and Waymo.
@@ 4.2.1 Perception Accuracy
Table 4: 3D semantic occupancy prediction comparison (%) on the validation set of Occ3D-nuScenes [33]. Sup. represents the supervised learning type. mIoU is the mean Intersection-over-Union excluding the ’others’ and ’other flat’ classes.
@@ 4.2.2 Inference Speed
Recent studies on 3D occupancy perception [76, 96] have begun to consider not only perception accuracy, but also its inference speed. According to the data provided by FastOcc [76] and FullySparse [96], we sort out the inference speeds of 3D occupancy methods, and also report their running platforms, input image sizes, backbone architectures, and occupancy accuracy on the Occ3D-nuScenes dataset, as depicted in Tab. 5.
@@ 5.1 Occupancy-based Applications in Autonomous Driving
3D occupancy perception enables a comprehensive understanding of the 3D world and supports various tasks in autonomous driving. Existing occupancy-based applications include segmentation, detection, flow prediction, and planning. (1) Segmentation: Semantic occupancy perception can essentially be regarded as a 3D semantic segmentation task. (2) Detection: OccupancyM3D [5] and SOGDet [133] are two occupancy-based works to implement 3D object detection. OccupancyM3D first learns occupancy to enhance 3D features, which are then used for 3D detection. SOGDet develops two concurrent tasks: semantic occupancy prediction and 3D object detection, training these tasks simultaneously for mutual enhancement. (3) Flow prediction: Cam4DOcc [9] predicts foreground flow in 3D space from the perspective of occupancy, and achieves an understanding of changes in the surrounding 3D environment. (4) Planning: OccNet [7] quantifies physical 3D scenes into semantic occupancy, and trains a shared occupancy descriptor. This descriptor is fed to various task heads to implement driving tasks. For example, the motion planning head outputs planning trajectories for ego vehicles.
However, existing occupancy-based applications primarily focus on the perception level, and less on the decision-making level. Given that 3D occupancy is more consistent with the 3D physical world than other perception manners (e.g., bird’s-eye view perception and perspective-view perception), we believe that 3D occupancy holds opportunities for broader applications in autonomous driving. At the perception level, it could improve the accuracy of existing trajectory prediction, 3D object tracking, and 3D lane line detection. At the decision-making level, it could help safer driving decisions and provide 3D explainability for driving behaviors.
@@ 5.3 Robust 3D Occupancy Perception
In dynamic and unpredictable real-world driving environments, the perception robustness is crucial to autonomous vehicle safety. State-of-the-art 3D occupancy models may be vulnerable to out-of-distribution scenes and data, such as changes in lighting and weather, which would introduce visual biases, and input image blurring, which is caused by vehicle movement. Moreover, sensor malfunctions (e.g., loss of frames and camera views) are common. In light of these challenges, studying robust 3D occupancy perception is valuable.
@@ 5.4 Generalized 3D Occupancy Perception
Furthermore, current 3D occupancy perception can only recognize a set of predefined object categories, which limits its generalizability and practicality. Recent advances in large language models (LLMs) [140, 141, 142, 143] and large visual-language models (LVLMs) [144, 145, 146, 147, 148] demonstrate a promising ability for reasoning and visual understanding. Integrating these pre-trained large models has been proven to enhance generalization for perception [8]. POP-3D [8] leverages a powerful pre-trained visual-language model [148] to train its network, and achieves open-vocabulary 3D occupancy perception. Therefore, we consider that employing LLMs and LVLMs is a challenge and opportunity for achieving generalized 3D occupancy perception.
@@ 6 Conclusion
This paper provided a comprehensive survey of 3D occupancy perception in autonomous driving in recent years. We reviewed and discussed in detail the state-of-the-art LiDAR-centric, vision-centric, and multi-modal perception solutions, and highlighted information fusion techniques for this field. To facilitate further research, detailed performance comparisons of existing occupancy methods are provided. Finally, we described some open challenges that could inspire future research directions in the coming years. We hope that this survey can benefit the community, support further development in autonomous driving, and help inexpert readers navigate the field.

### v2

@@ 3.3.1 General Pipeline
RGB images captured by cameras provide rich and dense semantic information but are sensitive to weather condition changes and lack precise geometric details. In contrast, point clouds from LiDAR or radar are robust to weather changes and excel at capturing scene geometry with accurate depth measurements. However, they only produce sparse features. Multi-modal occupancy perception can combine the advantages from multiple modalities, and mitigate the limitations of single-modal perception. Fig. 7 illustrates the general pipeline of multi-modal occupancy perception. Most multi-modal methods [11, 94, 12, 15] map 2D image features into 3D space and then fuse them with point cloud features. Moreover, incorporating 2D perspective-view features in the fusion process can further refine the representation [14]. The fused representation is processed by an optional refinement module and an occupancy head, such as 3D convolution or MLP, to generate the final 3D occupancy predictions. The optional refinement module [95] could be a combination of cross attention, self attention, and diffusion denoising [118].
@@ 3.4.1 Training with Strong Supervision
where $\left(V,\hat{V}\right)$ are the ground-truth and predicted semantic occupancy with $N_{C}$ categories. $\omega_{c}$ is a weight for a specific class $c$ according to the inverse of the class frequency. Moreover, some occupancy perception methods use other semantic losses, which are commonly utilized in semantic segmentation tasks [123, 124], such as Lovasz-Softmax loss [125] and Focal loss [126]. Furthermore, there are two specialized semantic losses: frustum proportion loss [25], which provides cues to alleviate occlusion ambiguities from the perspective of the visual frustum, and position awareness loss [127], which leverages local semantic entropy to encourage sharper semantic and geometric gradients.
@@ 3.4.2 Training with Other Supervisions
4 Evaluation
4.1 Datasets and Metrics
4.1.1 Datasets
There are a variety of datasets to evaluate the performance of occupancy prediction approaches, e.g., the widely used KITTI [129], nuScenes [86], and Waymo [130]. However, most of the datasets only contain 2D semantic segmentation annotations, which is not practical for the training or evaluation of 3D occupancy prediction approaches. To support the benchmarks for 3D occupancy perception, many new datasets such as Monoscene [25], Occ3D [35], and OpenScene [93] are developed based on the previous datasets like nuScenes and Waymo. A detailed summary of datasets is provided in Tab. 2.
@@ 4.2.1 Perception Accuracy
Table 4: 3D semantic occupancy prediction comparison (%) on the validation set of Occ3D-nuScenes [35]. Sup. represents the supervised learning type. mIoU is the mean Intersection-over-Union excluding the ’others’ and ’other flat’ classes. For fairness, all compared methods are vision-centric.
@@ 4.2.2 Inference Speed
Recent studies on 3D occupancy perception [82, 107] have begun to consider not only perception accuracy but also its inference speed. According to the data provided by FastOcc [82] and FullySparse [107], we sort out the inference speeds of 3D occupancy methods, and also report their running platforms, input image sizes, backbone architectures, and occupancy accuracy on the Occ3D-nuScenes dataset, as depicted in Tab. 5.
@@ 5.1 Occupancy-based Applications in Autonomous Driving
3D occupancy perception enables a comprehensive understanding of the 3D world and supports various tasks in autonomous driving. Existing occupancy-based applications include segmentation, detection, dynamic perception, and world models. (1) Segmentation: Semantic occupancy perception can essentially be regarded as a 3D semantic segmentation task. (2) Detection: OccupancyM3D [5] and SOGDet [144] are two occupancy-based works that implement 3D object detection. OccupancyM3D first learns occupancy to enhance 3D features, which are then used for 3D detection. SOGDet develops two concurrent tasks: semantic occupancy prediction and 3D object detection, training these tasks simultaneously for mutual enhancement. (3) Dynamic perception: Cam4DOcc [10] predicts foreground flow in 3D space from the perspective of occupancy and achieves an understanding of 3D environmental changes. (4) World models: OccNet [8] quantifies physical 3D scenes into semantic occupancy and trains a shared occupancy descriptor. This descriptor is fed to various task heads to implement driving tasks. For example, the motion planning head outputs planning trajectories for ego vehicles, and the detection head outputs 3D boxes of objects. Similarly, UniScene [60], Occupancy-MAE [99] and DriveWorld [7] pre-train occupancy to extract a universal representation of the 3D scene, and then fine-tune various downstream tasks, such as 3D object detection, online mapping, multi-object tracking, motion forecasting, occupancy prediction, and planning.
However, existing occupancy-based applications primarily focus on the perception level, and less on the decision-making level. Given that 3D occupancy is more consistent with the 3D physical world than other perception manners (e.g., bird’s-eye view perception and perspective-view perception), we believe that 3D occupancy holds opportunities for broader applications in autonomous driving. At the perception level, it could improve the accuracy of existing place recognition [145], pedestrian detection [146, 147], accident prediction [148], and lane line segmentation [149]. At the decision-making level, it could help safer driving decisions [150] and navigation [151, 152], and provide 3D explainability for driving behaviors.
@@ 5.3 Robust 3D Occupancy Perception
In dynamic and unpredictable real-world driving environments, the perception robustness is crucial to autonomous vehicle safety. State-of-the-art 3D occupancy models may be vulnerable to out-of-distribution scenes and data, such as changes in lighting and weather, which would introduce visual biases, and input image blurring, which is caused by vehicle movement. Moreover, sensor malfunctions (e.g., loss of frames and camera views) are common [153]. In light of these challenges, studying robust 3D occupancy perception is valuable.
@@ 5.4 Generalized 3D Occupancy Perception
Furthermore, current 3D occupancy perception can only recognize a set of predefined object categories, which limits its generalizability and practicality. Recent advances in large language models (LLMs) [160, 161, 162, 163] and large visual-language models (LVLMs) [164, 165, 166, 167, 168] demonstrate a promising ability for reasoning and visual understanding. Integrating these pre-trained large models has been proven to enhance generalization for perception [9]. POP-3D [9] leverages a powerful pre-trained visual-language model [168] to train its network and achieves open-vocabulary 3D occupancy perception. Therefore, we consider that employing LLMs and LVLMs is a challenge and opportunity for achieving generalized 3D occupancy perception.
@@ 6 Conclusion
This paper provided a comprehensive survey of 3D occupancy perception in autonomous driving in recent years. We reviewed and discussed in detail the state-of-the-art LiDAR-centric, vision-centric, and multi-modal perception solutions and highlighted information fusion techniques for this field. To facilitate further research, detailed performance comparisons of existing occupancy methods are provided. Finally, we described some open challenges that could inspire future research directions in the coming years. We hope that this survey can benefit the community, support further development in autonomous driving, and help inexpert readers navigate the field.
Acknowledgment
The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust.